# T-3DGS: Removing Transient Objects for 3D Scene Reconstruction

### v1

@@ E Handling Semi-Transient Objects
Semi-transient objects are have not been properly addressed in 3D scene reconstruction. Our method makes a large improvement over previous work and can handle fairly complex scenarios. We provide some details about data capture and the methodology we employed with respect to semi-transient objects. We also discuss why both our consistency term as well as mask propagation are important for handling semi-transient objects. Additionally, we discuss some limitations of our proposed method.
@@
Some scenes in our dataset include reduced and full versions - in the reduced scenes the camera operator goes from one end of the scene to the other, while for the full scenes they also retrace their path and return to their initial positions, while semi-transient objects kept moving. As seen in Fig. 10, our proposed TMR module is crucial to produce good results even for the reduced scenes, the more challenging scenario. For full scenes, the additional data significantly improves mask predictions for all models. We conclude that handling semi-transient objects is an important challenge for in-the-wild video processing. In order to provide the most challenging datasets and to rigorously compare different methods - one must consider not only the kinds of motion of dynamic objects, but also their movement with respect to the camera.
@@
As mentioned above, some of the scenes include semi-transient objects occluding the static scene for prolonged periods of time while remaining mostly still. As this period of time increases, semi-transient objects can effectively become static. While this effect might seem irrelevant to the detection of dynamic objects, this is not the case. As shown in the Fig. 11, most methods mask static background as if it was masking the semi-transient object. Notably, our method avoids this problem, and this highlights the importance of using both the consistency term and mask propagation algorithm we introduced. While the consistency term alone may still misclassify semi-transient objects, we avoid false classifications of static objects as dynamic. Therefore, our consistency term is crucial in order for the mask propagation to avoid introducing additional errors. This is in stark contrast to the competing methods, which have a lot more false positives. Mask propagation would introduce additional errors and may not help improving the overall quality.
@@
Our method reliably removes transient and semi-transient distractors and successfully reconstructs static artifact-free 3D scenes. However, we have noticed that predicted masks are inflated due to the low resolution of the extracted feature maps, our method can give inconsistent results for small objects, since DINOv2 features are computed on patches. These problems could be addressed by utilizing feature extractors with higher-resolution feature maps or guided upsampling. Additionally, the temporal refinement process is constrained by a memory window of $N_{m}$ frames, which means that if an object disappears for more than $N_{m}$ frames and then reappears, it will be treated as a new instance with a different label. This can lead to inconsistent tracking and potentially affect the filtering process, especially for semi-transient objects that may temporarily leave the scene. Furthermore, our current filtering approach using global coverage scores may incorrectly filter out valid dynamic objects that undergo significant size changes, such as objects moving towards or away from the camera, or those experiencing perspective changes. We leave it as future work.

### v2

@@ Appendix F Handling Semi-Transient Objects
Semi-transient objects have not been properly addressed in 3D scene reconstruction. Our method represents a significant improvement over previous work and can handle relatively complex scenarios. We provide details on the data capture process and the methodology employed for handling semi-transient objects. We also discuss the importance of both divergence estimation and mask propagation in handling semi-transient objects. Additionally, we discuss the limitations of our proposed method.
@@
Our dataset includes two versions of some scenes: reduced and full. In reduced scenes, the camera operator moves from one end of the scene to the other. In full scenes, however, the operator retraces this path back to the starting position while semi-transient objects continue to move. As illustrated in Fig. 8, our proposed TMR module is essential for achieving good results in reduced scenes, which are particularly challenging. In full scenes, the additional frames lead to significantly improved mask predictions for all models because transient objects remain visible for longer periods. When fewer frames capture the scene, many methods mistakenly classify these transient objects as static. Overall, our findings highlight that effectively handling semi-transient objects is a major challenge in in-the-wild video processing. To develop the most challenging datasets and to rigorously compare different methods, it is important to consider not only the types of motion dynamic objects exhibit but also their movement relative to the camera.
@@
As mentioned above, some of the scenes include semi-transient objects occluding the static scene for prolonged periods of time while remaining mostly still. As this period of time increases, semi-transient objects can effectively become static. Although this effect might seem irrelevant to the detection of dynamic objects, this is not the case. As shown in the Fig. 9, most methods mask the static background as if it were masking the semi-transient object. Notably, because WildGaussians relies heavily on semantic information, it can ”propagate” the masks. However, this happens too late into the training process while our method avoids this problem, and this highlights the importance of using both divergence estimation and mask propagation algorithm we have proposed. Moreover, we aim to minimize false classifications of static objects as dynamic. As discussed earlier, even WildGaussians produces an excessive number of misclassifications for TMR. Therefore, our method is crucial for mask propagation to avoid introducing additional errors. This is in stark contrast to the competing methods, which have a lot more false positives. Mask propagation could introduce additional errors and might not contribute to overall quality improvement.
@@
Our method reliably removes transient and semi-transient distractors and successfully reconstructs static artifact-free 3D scenes. However, we have observed that predicted masks tend to be inflated due to the low resolution of the extracted feature maps. Our method can also produce inconsistent results for small objects, as DINOv2 features are computed on patches. These problems could be addressed by using feature extractors with higher-resolution feature maps or guided upsampling. Additionally, the temporal refinement process is limited by a memory window of $N_{m}$ frames, which means that if an object disappears for more than $N_{m}$ frames and then reappears, it will be treated as a new instance with a different label. This can lead to inconsistent tracking and potentially affect the filtering process, especially for semi-transient objects that may temporarily leave the scene. Furthermore, our current filtering approach using global coverage scores may incorrectly filter out valid dynamic objects that undergo significant size changes, such as objects moving towards or away from the camera, or those experiencing perspective changes. We leave this aspect for future work.